OpenStudy (anonymous):

Please help! What is the remainder of (x^3 – 6x – 9x + 3) ÷ (x - 3) ? A. -51 B. -51/x - 3 C. -17/x D. -17/x - 3 A, B, C, or D? Please explain (:

OpenStudy (raden):

i think it will easier if use the "remainder theorema", f(x) : (x-a), so its remainder is f(a) now, if f(x) = x^3 – 6x – 9x + 3 then f(3) = .... ?

OpenStudy (anonymous):

I have no idea.

OpenStudy (raden):

just plug x=3, u will get f(3)

OpenStudy (raden):

f(x) = x^3 – 6x – 9x + 3 f(3) = 3^3 -6(3) - 9(3) + 3 = ....

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

Yes?

OpenStudy (raden):

ok... correct

OpenStudy (raden):

but, no option there :)

OpenStudy (anonymous):

Right, so. Any idea on what the answer might be?

OpenStudy (raden):

i think there is a typo error above, maybe... is ur question : x^3 – 6x – 9x + 3 or x^3 – 6x^2 – 9x + 3 ?

OpenStudy (raden):

yea, the second term should be -6x^2, right ?

OpenStudy (anonymous):

No. Just 6x is the second term.

OpenStudy (raden):

actually, a polynomial has a ascending order degree : f(x) = ax^n +- ax^(n-1) +- ax^(n-2) +- ax^(n-3) +- k why, the polynomial above be x^3 – 6x – 9x + 3, because it can be simplified to x^3 - 15x + 3, right ?

OpenStudy (raden):

i mean, f(x) = ax^n +- ax^(n-1) +- ax^(n-2) +- ax^(n-3) +- ..... +- k +-, means can be + or -

OpenStudy (raden):

if ur question is f(x) = x^3 – 6x^2 – 9x + 3, so f(3) : f(3) = 3^3 -6(3)^2 -9(3) + 3 = .... u will get the answer in option :)

OpenStudy (anonymous):

Ugh. I'm never gonna find out this freaking answer.